The Variable Speed Company manufactures a line of high quality tools. The company sold 1,050,000 hammers at a price of $4.5 per unit last year. The company estimates that this volume represents a 20% share of the current hammers market. The market is expected to increase by 6%. Marketing specialists have determined that, as a result of a new advertising campaign and packaging, the company will increase its share of this larger market to 24%. Due to changes in prices, the new price for the hammer will be $4.80 per unit. This new price is expected to be in line with the competition and have no effect on the volume estimates. What are the estimated sales revenues in the coming year?

Multiple Choice

  • $6,048,000.

  • $6,010,200.

  • $6,795,633.

  • $6,410,880.
Answer #1

Current market hammers = 1,050,000 / 20% = 5,250,000 hammers

Expected market hammers = 5,250,000 + 6% = 5,565,000

Estimated sales revenue = 5,565,000 * 24% * $4.80 = $6,410,880